OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 8 − 5x = 4x − 1 4x + 4 = 10 − 2x The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describes solutions to equations shown.

OpenStudy (anonymous):

The solution to equation #1 is the same as the solution to equation #2.

OpenStudy (ash2326):

Ghass would you please explain how you got this answer?

OpenStudy (anonymous):

eqn1 :8 − 5x = 4x − 1 9=9x x=1 eqn2. 4x + 4 = 10 − 2x 6x=6 x=1 Both are x=1 Therefore, The solution to equation #1 is the same as the solution to equation #2.